3w-13=1/4(52-12w)
3w-13=13-3w


First, i know i have to distribute 1/4 to 52 and -12w.
Then, i'm stuck, i know that the answer is 4 1/3 but i don't know how to get that answer. PLEASE HELPP!

Answer by stanbon(48568)

3w-13=1/4(52-12w)
3w-13=13-3w
----------------
This latter equation is true for all values of w;
it is an identity.
-------------------
The solution is all Real Numbers.
Cheers,
Stan H.